# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v1

@@ 5.6 Predictive monitoring
In modern industry, sensor-based monitoring of processes has become essential. This pairs with the rapid growth of the Internet of Things (IoT), which aims at a smarter management across various applications [art_9]. In this framework, multi-sensor systems can be used to detect anomalies in complex scenarios and to monitor the overall status of a system. Predictive monitoring is a task that involves the continuous observation and analysis of a system, in order to forecast its future states and verify whether the predicted outcomes meet certain standards. This predictive monitoring task is important because it significantly improves efficiency, reliability, and cost of industrial operations. Related approaches include several techniques, such as anomaly detection, fault diagnosis and remaining useful life estimation.
Despite the specificity of the field, 16 out of 156 papers address the topic of predictive monitoring. This includes anomaly detection ([art_9], [art_10]), fault diagnosis ([art_53], [art_54], [art_55], [art_56], [art_57], [art_58], [art_192]) and RUL estimation ([art_152], [art_153], [art_154], [art_155], [art_156], [art_157], [art_158]). The use of GNNs in these analyses is justified by the necessity of using a model capable of capturing the spatio-temporal correlations among the many sensors present in the system.
@@ 5.6.1 Overview
Predictive monitoring includes several techniques, such as anomaly detection, fault diagnosis and remaining useful life estimation. Anomaly detection is a technique that aims to identify abnormal data that are not due to random deviations, and are rather generated by a different underlying mechanism [ref_anomaly_detection_1]. It is a prerequisite part of fault diagnosis [ref_anomaly_detection_2]. Fault diagnosis is the process of determining whether a fault has occurred in a system, including the identification of the time, location, type, and severity of the fault. It is usually considered to be a classification problem. As for remaining useful life (RUL) estimation, it aims to determine how long a machine will operate before it needs to be repaired or replaced, and it is useful for scheduling maintenance interventions.
The selected papers implement these three techniques, with anomaly detection being less prevalent than the other two. The main applications studied are industrial (e.g., bearings fault diagnosis [art_54], [art_55], [art_57], [art_155], [art_156]) and mechanical (e.g, engines RUL estimation [art_152], [art_153], [art_154], [art_157]). There are also two papers dealing with energy related tasks, specifically fault diagnosis of photovoltaic systems [art_53] and energy networks [art_56]. They are included in this section due to their specific focus on fault classification rather than on energy-related time series forecasting. It shoud also be underlined that the majority of the papers aim at solving a classification task, which is typical of anomaly detection and fault diagnosis. As for RUL estimation, the quantity to be predicted is the RUL itself. For this reason, we do not refer to forecasting time horizons as in the other sections, since the goal is to estimate the RUL rather than to forecast data over future time intervals.
The majority of the collected papers were published in 2023 by authors affiliated with Chinese universities. The most popular journal for these topics is IEEE Transactions on Instrumentation and Measurement by the Institute of Electrical and Electronics Engineers by IEEE, with 5 out of 15 papers. All other journals contain only one of the selected articles. It can be observed that the majority of the articles are published in IEEE journals.
@@ 5.6.2 Datasets
Table 25: List of public datasets in the ”Predictive monitoring” group and their corresponding links.
@@
C-MAPSS [art_152], [art_153], [art_154], [art_157] https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf
CWRU bearing dataset [art_55], [art_192] http://csegroups.case.edu/bearingdatacenter/home
@@
XJTUGearbox and XJTUSpurgear [art_192] http://dx.doi.org/10.1016/j.ymssp.2021.108653 (source paper)
@@
The two most popular datasets are C-MAPSS and N-CMAPSS, two well-known datasets for RUL prediction provided by NASA. C-MAPSS describes the deterioration of aircraft engines by recording temperature, pressure, and fan speed during the cruise phase (https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf). N-CMAPSS is an extended version of C-MAPSS, whose records cover climbing, cruising, and descending flight conditions (https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/). Practically no papers use exogenous variables, with the exception of Ref. [art_53], which evaluates the impact of meteorological variables on fault diagnosis of photovoltaic systems.
@@
The granularity of the studied time series is highly variable, ranging from 1 hour to fractions of seconds. Many papers report the frequency of the sensors, with values ranging from 1 Hz to 50 kHz. Most datasets are normalized, by either using min-max normalization or Z-score normalization. In some papers, a data augmentation procedure is employed to mitigate the natural imbalance between the number of normal and anomalous conditions.
5.6.3 Proposed models
@@
Half of the proposed GNN models (9 out of 16 papers) fall under the category of pure convolutional models, followed by 4 attentional GNNs, 2 hybrid models, and one whose description does not clarify which class of the taxonomy it belongs to (Ref. [art_155]). Almost all the models have a graph learning module, so that the graph structure is rarely defined a priori by the researchers. Three papers ([art_55], [art_57], [art_154]) use the visibility graph algorithm proposed by Lacasa et al. in Ref. [visibility_graph] for converting the time series into a graph.
The most common loss functions in this group are the cross entropy function for classification problems, and MSE or RMSE for RUL forecasting problems. Most of the papers mention the Python language and the PyTorch library, but only two papers provide a link to the source code of their model, and in both cases, it is for RUL prediction. The source code for the convolution-graph attention network (ConvGAT) [art_153] is available at https://github.com/CUG-FDGroup, while the one for the graph network with causal connectivity and temporal convolutional neural network feature extractor proposed in Ref. [art_155] is available at https://github.com/mylonasc/gnn-tcnn.
@@ 5.6.4 Benchmark models
Tab. LABEL:t:benchmark_models_predictive_monitoring presents the benchmark models employed in the ”Predictive monitoring” group by at least two articles examined in this review.
Table 26: List of benchmark models in the ”Predictive monitoring” group divided per category.
@@
Classical machine learning methods Attention-based long-short term memory network [art_153], [art_157]
Feature-attention based bidirectional GRU and CNN model (AGCNN) [AGCNN] [art_153], [art_154], [art_157]
Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_152], [art_153], [art_154], [art_157]
Convolutional neural network (CNN) [art_9], [art_56], [art_152], [art_153], [art_154], [art_156], [art_157]
Decision tree (DT) [art_9], [art_56]
Feed-forward neural network (FNN) [art_55], [art_155]
Gradient boosting (GB) [art_56], [art_154]
Gated recurrent unit (GRU) [art_53], [art_56], [art_156]
@@
Long-short term memory network (LSTM) [art_55], [art_56], [art_154], [art_157], [art_158]
Random forest (RF) [art_56], [art_154]
Support vector machine (SVM) [art_9], [art_56]
Transformer [attention] [art_152], [art_154]
GNN methods Graph attention network (GAT) [GAT] [art_54], [art_56], [art_192]
Graph convolutional network (GCN) [art_54], [art_55], [art_156], [art_192]
Hierarchical attention graph convolutional network (HAGCN) [HAGCN] [art_152], [art_153]
Spatio-temporal fusion attention (STFA) [art_157] [art_153], [art_154]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_58], [art_152]
In this group of papers there is a high variety of benchmarks used, and only few papers share a benchmark model, especially in the GNN category. The most popular benchmarks are two classic well-known machine learning models, the CNN and the LSTM network. Not surprisingly, there are no classical statistical or mathematical benchmark models, since they are more frequently used in forecasting problems compared to classification problems.
Among the GNN benchmarks, there are two models that have been specifically developed for remaining useful life prediction, namely the hierarchical attention graph convolutional network (HAGCN) [HAGCN] and the spatio-temporal fusion attention (STFA) [art_157]. In the HAGCN model, proposed by Li et al. in 2021, the spatial dependencies are modeled by a hierarchical graph representation layer, and a bidirectional long-short term memory network is used for modeling temporal dependencies of sensor measurements. As for the STFA model, proposed by Kong et al. in 2022, it integrates a priori knowledge about the equipment’s structure with a spatio-temporal deep learning architecture which employs LSTM cells and an attention mechanism. Neither of the two original papers includes a link to the source code.
@@ 5.6.5 Results
The results demonstrate that the proposed GNN models outperform the benchmarks. The accuracy of the models is evaluated using different metrics, depending on whether the context is classification or forecasting. For classification tasks, several studies present confusion matrices and calculate accuracy percentages. For RUL forecasting problems instead, the most commonly used metrics are the RMSE and the score function.
@@
Table 27: Comparison of the accuracy of the different models on the C-MAPSS dataset on sub-datasets FD001–FD004, expressed in terms of RMSE and score function. The smallest errors are underlined. Numbers from the papers.
@@
Transformer 13.52 287.07 19.32 1436.74 13.44 263.64 20.38 2784.62
In general, the best results on the different sub-datasets are achieved by the GNN-based comprehensive dynamic structure GNN (CDSG) [art_154] and convolution-graph attention network (ConvGAT) [art_153] models.
@@ 5.7 Generic
This subsection includes all the papers that do not explicitly address a specific problem. They still present the results on some benchmark case studies, but the related datasets may cover different groups. In particular, the focus is on papers that do not discuss or mention any specific case study prior to the experimental section. The reason for this choice is that they are considered to explore broader methodologies or frameworks that are applied to diverse datasets without a declared focus on a particular case.
@@ 5.7.1 Overview
A significant number of papers (24 out of 156) have been categorized under the ”Generic” group. The majority of these works commit to address a multivariate time series forecasting problem which is not necessarily limited to a specific field. This ”Generic” subset includes one paper from 2021, 4 from 2022, and 19 from 2023, highlighting the recent interest in applying GNNs in more general contexts. These papers are distributed across various journals, with the exception of Expert Systems with Applications by Elsevier, IEEE Transactions on Knowledge and Data Engineering by IEEE, and Information Sciences by Elsevier, each publishing 2 of the selected papers. The majority of authors are affiliated with institutions in China.

### v2

@@ 5.6 Predictive monitoring
The theme of predictive monitoring is related to sensor-based monitoring of processes, which has become essential in modern industry. The evolution of monitoring systems is concomitant with the rapid growth of the Internet of Things (IoT), which aims at a smarter management across various applications [art_9]. Hence, in this context, multi-sensor systems can be used to detect anomalies in complex scenarios and to monitor the overall status of a system. In the specific, predictive monitoring is a task that involves the continuous observation and analysis of a system in order to forecast its future states and verify whether the predicted outcomes meet certain standards. The predictive monitoring task is important because it can significantly improve efficiency, reliability, and cost of industrial operations. Related data-based approaches include techniques such as anomaly detection, fault diagnosis, and estimation of the equipment’s remaining useful life.
Despite the specificity of the field, 35 out of 281 papers address the topic of predictive monitoring. This includes anomaly detection [art_9, art_10, art_477, art_479, art_481, art_483, art_484, art_489], fault diagnosis [art_53, art_54, art_55, art_56, art_57, art_58, art_192, art_469, art_471, art_474, art_486, art_487, art_491] and RUL estimation [art_152, art_153, art_154, art_155, art_156, art_157, art_158, art_478, art_488]. The use of GNNs in these analyses is justified by the necessity of using a model capable of capturing the spatio-temporal correlations among the many sensors present in the system.
@@ 5.6.1 Overview
As said, predictive monitoring is a research field which focuses on several techniques, such as anomaly detection, fault diagnosis and remaining useful life estimation. Anomaly detection is a technique that aims to identify abnormal data that are not due to random deviations from some regular pattern, but are generated by a different underlying mechanism [ref_anomaly_detection_1]. It is a prerequisite part of fault diagnosis [ref_anomaly_detection_2]. Fault diagnosis is the process of determining whether a fault has occurred in a machine, including the identification of the time, location, type, and severity of the fault. It is usually considered to be a classification problem. As for remaining useful life (RUL) estimation, it aims to determine how long a machine will operate before it needs to be repaired or replaced, and it is useful for scheduling maintenance interventions.
The selected papers implement these three techniques, with anomaly detection being less studied than the other two. The main applications developed are industrial (e.g., bearings fault diagnosis [art_54, art_55, art_57, art_155, art_156]) and mechanical (e.g, engines RUL estimation [art_152, art_153, art_154, art_157, art_478, art_488]). There are also some papers dealing with energy related tasks, specifically fault diagnosis of photovoltaic systems [art_53] and of energy networks [art_56, art_487]. They are included in this thematic group because of their specific focus on fault classification rather than on energy-related time series forecasting. It should also be underlined that the majority of these papers aim at solving a classification task, which is typical of anomaly detection and fault diagnosis. As for RUL estimation, the quantity to be predicted is the RUL itself. For this reason, we do not make specific reference to forecasting time horizons, unlike in the other subsections, since the goal here is to estimate the RUL rather than to forecast data over future time intervals.
The majority of the collected papers related to this theme were published in 2024 by authors affiliated with Chinese universities. The most popular journal for these topics is IEEE Transactions on Instrumentation and Measurement by IEEE, with 6 out of 35 papers. It can be observed that the majority of the articles are published in IEEE journals.
@@ 5.6.2 Datasets
Table 29: List of public datasets in the “Predictive monitoring” group and their corresponding links.
@@
C-MAPSS [art_152], [art_153], [art_154], [art_157], [art_478], [art_488] https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf
CWRU bearing dataset [art_55], [art_192], [art_491] http://csegroups.case.edu/bearingdatacenter/home
Herbert river water level [art_483] https://zenodo.org/records/8053359
Intelligent maintenance system (IMS) [art_475] https://www.nasa.gov/intelligent-systems-division
@@
Mars science laboratory rover (MSL) [art_481], [art_489] https://doi.org/10.1145/3219819.3219845 (source paper)
@@
Pooled server metrics (PSM) [art_479], [art_481] https://doi.org/10.1145/3447548.3467174 (source paper)
Server Machine Dataset (SMD) [art_479], [art_484] https://github.com/NetManAIOps/OmniAnomaly
@@
Secure water treatment (SWaT) [art_477], [art_479], [art_481], [art_484], [art_489] https://itrust.sutd.edu.sg/itrust-labs_datasets/dataset_info/
Soil moisture active passive satellite (SMAP) [art_479], [art_481], [art_484], [art_489] https://doi.org/10.1145/3219819.3219845 (source paper)
Water distribution (WaDi) [art_477], [art_479], [art_484], [art_489] https://itrust.sutd.edu.sg/itrust-labs_datasets/dataset_info/
Wireless capsule endoscopy images [art_475] https://doi.org/10.1055/s-0043-105488 (source paper)
XJTUGearbox and XJTUSpurgear [art_192], [art_474] http://dx.doi.org/10.1016/j.ymssp.2021.108653 (source paper)
@@
The two popular datasets are C-MAPSS and N-CMAPSS, two well-known datasets for RUL prediction provided by NASA. C-MAPSS describes the deterioration of aircraft engines by recording temperature, pressure, and fan speed during the cruise phase (https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf). N-CMAPSS is an extended version of C-MAPSS, whose records cover climbing, cruising, and descending flight conditions (https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/). Almost no papers use exogenous variables, with the exception of Ref. [art_53], which evaluates the impact of meteorological variables on fault diagnosis of photovoltaic systems.
@@
The time granularity of the studied time series is highly variable, ranging from 1 hour to fractions of seconds. Many papers report the frequency of the studied sensors, with values ranging from 1 Hz to 50 kHz. Most datasets are normalized, by either using min-max normalization or Z-score normalization. In some papers, a data augmentation procedure is employed to mitigate the natural imbalance between the number of normal and anomalous conditions.
5.6.3 Types of models
@@
More than half of the proposed spatio-temporal GNN models (19 out of 35 papers) fall under the category of pure convolutional GNNs, followed by 10 attentional GNNs, 2 hybrid models, and one recurrent model (Ref. [art_470]). One paper (Ref. [art_57]) uses both convolutional and attentional GNNs, one paper (Ref. [art_475]) uses a variant of the convolutional GNN, and the description of another one (Ref. [art_155]) does not clarify which class of the taxonomy it belongs to. Almost all the models have a graph learning module, so that the graph structure is rarely defined a priori by the researchers. Three papers [art_55, art_57, art_154] use the visibility graph algorithm proposed by Lacasa et al. in Ref. [visibility_graph] for converting the time series into a graph.
The most common loss functions used within this thematic group are the cross entropy function for classification problems, and MSE or RMSE for RUL forecasting problems. Most of the papers mention the Python language and the PyTorch library, but only the papers listed in Tab. LABEL:t:link_codici_predictive_monitoring provide the links to the related source codes.
Table 30: List of source codes of the “Health” models in the review.
Model Link
Convolution-graph attention network (ConvGAT) [art_153] https://github.com/CUG-FDGroup
Graph Network with Causal Connectivity and Temporal Convolutional Neural Network feature extractor [art_155] https://github.com/mylonasc/gnn-tcnn
Graph spatiotemporal process and anomaly scorer (GST-Pro) [art_477] https://github.com/huankoh/GST-Pro
Causality Enhanced Global-Local Graph Neural Network (CEGLo-GNN) [art_482] https://github.com/YueYueXia/CEGLo-GNN
Graph Deviation Network+ (GDN+) [art_483] https://github.com/KatieBuc/gnnad/releases
@@ 5.6.4 Benchmark models
Tab. LABEL:t:benchmark_models_predictive_monitoring presents the benchmark models employed in the “Predictive monitoring” thematic group by at least two articles examined.
Table 31: List of benchmark models in the “Predictive monitoring” group divided per category.
@@
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_483], [art_487]
Non-GNN machine learning methods Attention-based long-short term memory network [art_153], [art_157], [art_488]
Feature-attention based bidirectional GRU and CNN model (AGCNN) [AGCNN] [art_153], [art_154], [art_157], [art_478], [art_488]
Anomalous Rhythm Detection using Adversarially Generated Time Series (BeatGAN) [BeatGAN] [art_481], [art_484]
Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_152], [art_153], [art_154], [art_157], [art_470], [art_488]
Convolutional neural network (CNN) [art_9], [art_56], [art_152], [art_153], [art_154], [art_156], [art_157], [art_461], [art_469], [art_471], [art_486], [art_487], [art_488]
Convolutional neural network with long-short term memory network (CNN-LSTM) [art_154], [art_470]
Crossformer [Crossformer] [art_152], [art_482]
Deep autoencoding gaussian model (DAGMM) [DAGMM] [art_481], [art_484]
Deep belief network (DBN) [DBN] [art_157], [art_488]
Decision tree (DT) [art_9], [art_56], [art_487]
Feed-forward neural network (FNN) [art_55], [art_155], [art_469], [art_478]
Genetic algorithm approach on restricted Boltzmann-based long-short term memory network (GA+RBM+LSTM) [GA+RBM+LSTM] [art_157], [art_488]
Gradient boosting (GB) [art_56], [art_154], [art_487]
Gated recurrent unit (GRU) [art_53], [art_56], [art_156], [art_469], [art_487]
Integrated multi-head dual sparse self-attention network (IMDSSN) [IMDSSN] [art_154], [art_478], [art_488]
Isolation forest [art_10], [art_481]
$k$-nearest neighbors (KNN) [art_56], [art_475], [art_487]
Local outlier factor (LOF) [art_475], [art_481]
@@
Long-short term memory network (LSTM) [art_55], [art_56], [art_154], [art_157], [art_158], [art_461], [art_470], [art_487], [art_488]
Long-short term memory-based variational auto-encoder (LSTM-VAE) [art_10], [art_477], [art_479], [art_484], [art_489]
OmniAnomaly [OmniAnomaly] [art_477], [art_479], [art_481], [art_484], [art_489]
Residual network (ResNet) [ResNet] [art_471], [art_491]
Random forest (RF) [art_56], [art_154], [art_478], [art_487]
Support vector machine (SVM) [art_9], [art_56], [art_471], [art_478], [art_486], [art_487]
Deep transformer networks for anomaly detection (TranAD) [TranAD] [art_484], [art_489]
Transformer [attention] [art_152], [art_154], [art_470], [art_478], [art_479], [art_488], [art_489]
Unsupervised anomaly detection on multivariate time series (USAD) [5] [art_479], [art_484], [art_489]
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_58], [art_482]
Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_480], [art_482]
Fused sparse autoencoder and graph network (FuSAGNet) [FuSAGNet] [art_10], [art_479]
Graph attention network (GAT) [GAT] [art_54], [art_56], [art_192], [art_470], [art_487]
Graph convolutional network (GCN) [art_54], [art_55], [art_156], [art_192], [art_461], [art_474], [art_486]
Graph deviation network (GDN) [GDN] [art_10], [art_477], [art_479], [art_481], [art_484]
Gated graph convolutional network (GGCN) [GGCN] [art_153], [art_478]
Graph WaveNet (GWN) [GWN] [art_58], [art_480], [art_482]
Hierarchical attention graph convolutional network (HAGCN) [HAGCN] [art_152], [art_153], [art_478]
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_470], [art_483]
Spatio-temporal fusion attention (STFA) [art_157] [art_153], [art_154], [art_478], [art_488]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_58], [art_152], [art_470], [art_478], [art_480]
In this thematic group there is a high variety of benchmarks used, and before 2024 only few papers share a benchmark model in the GNN category. The most popular benchmarks are two classic machine learning models, the CNN and the LSTM network. Not surprisingly, almost no statistical or mathematical benchmark models were used, since they are more frequently used in forecasting problems than in classification problems.
Among the GNN benchmarks, some of them have been specifically developed for remaining useful life prediction or predictive monitoring tasks, such as the hierarchical attention graph convolutional network (HAGCN) [HAGCN] and the spatio-temporal fusion attention (STFA) [art_157]. In the HAGCN model, proposed by Li et al. in 2021, the spatial dependencies are modeled by a hierarchical graph representation layer, and a bidirectional long-short term memory network is used for modeling temporal dependencies of sensor measurements. As for the STFA model, proposed by Kong et al. in 2022, it integrates some a priori knowledge about the equipment’s structure with a spatio-temporal deep learning architecture which employs LSTM cells and an attention mechanism. Neither of the two original papers includes a link to the source code.
@@ 5.6.5 Results
The reported results show that the proposed GNN models clearly outperform the benchmarks. The accuracy of the models is evaluated using different metrics, depending on whether the context is classification or forecasting. For classification tasks, several studies present confusion matrices and calculate accuracy percentages. For RUL forecasting problems instead, the most commonly used metrics are the RMSE and the score function.
@@
Table 32: Comparison of the accuracy of the different models on the C-MAPSS dataset on sub-datasets FD001–FD004, expressed in terms of RMSE and score function. The smallest errors are underlined. Numbers from the original papers.
@@
STP-GNN [art_488] 11.38 192.17 15.81 1094.81 11.02 199.58 17.77 1692.19
TATFA-Transformer 12.21 261.50 15.07 1359.70 11.23 210.21 18.81 2506.35
Transformer 12.92 276.27 22.64 1960.25 13.04 284.41 24.81 3892.24
In general, the best results on the different sub-datasets are achieved by the convolution-graph attention network (ConvGAT) [art_153] and the the dual-stream spatio-temporal fusion network (DS-STFN) [art_478] models. The ConvGAT model, proposed by Chen et al. in 2023, applies independent convolutional kernels to each feature of the input time series, and passes the resulting values to a graph attention network, with graph structure defined using cosine similarity. Finally, a fully connected layer outputs the RUL prediction. The DS-STFN model, proposed by Zhang et al. in 2024, organizes the data into grid-structured and graph-structured formats, with the graph structure defined by cosine similarity. A temporal convolutional network with self-attention extracts temporal features from the grid data, and, in parallel, a convolutional graph transformer module captures spatial features from the graph data. The spatial and temporal features are then fused to predict the RUL.
@@ 5.7 Generic
This subsection includes all the papers that do not specifically address a problem in the specific themes discussed so far. These papers present results on benchmark case studies, for which the related datasets span different thematic groups at once. In particular, the focus is on papers that do not discuss or mention any specific case study prior to the experimental section, or analyze datasets from multiple fields. The reason for this choice is that they are considered to explore broader methodologies or frameworks that are applied to diverse datasets without a declared focus on a particular case.
@@ 5.7.1 Overview
A significant number of papers (44 out of 281) have been categorized under the “Generic” thematic group. The majority of these works commit themselves to address a multivariate time series forecasting problem not necessarily limited to a specific field. This “Generic” set includes one paper from 2021, 4 from 2022, 19 from 2023, and 20 from 2024, highlighting the recent interest in applying GNNs in cross-disciplinary contexts. These papers are scattered across various journals, with the exception of Expert Systems with Applications by Elsevier, IEEE Transactions on Knowledge and Data Engineering by IEEE, and Information Sciences by Elsevier, which published 4 and 3 of the selected papers, respectively. The majority of authors are affiliated with institutions in China.